Determine the force of gravitational attraction between the earth (m=5.98×10²⁴kg)and a 70-kg STEM student if the students is standing at sea level,a distance of 6.38× 10^6 m from earth's center.

1
Expert's answer
2021-11-23T10:31:34-0500
"F_g=\dfrac{GMm}{R^2},""F_g=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times5.98\times10^{24}\ kg\times70\ kg}{(6.38\times10^6\ m)^2},""F_g=686\ N."